• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name perdrisat

    Log entry time 00:12:47 on June 2,2002

    Entry number 85054

    keyword=shift summary (swing)

    continue the current scan on LD2
    start run 1071 at 16:00, 70 mA
    80 mA starts at 1072, MPMA off, no good
    1074 still bad beam position
    1075 Coda crash, 400 k on tape
    1076 Coda problems, called Bob Michaels, who helped us out over the phone
    1077 the good run at 80 mA
    1078 go to 90 mA at 17:20
    1079 at 94 mA, maximum that can be achieved today (would require spot change)
    18:00 change tgt to C thick; machine tripped because forgot to request tgt masking.
    no problem! should never happen again.
    1080 C with 94 mA, raster on, 240 k events
    1081 same without raster, 92 mA! keep raster off from now on.
    1082 80 mA; thunder hits 18:25, machine down, 9000 events
    19:10 accelerator restarts, we are promised 100 mA, ask for 110, told 105 is hard limit (rad. mon.)
    So we go back to 80 mA.
    1083 80 mA 550 kevents
    1084 70 mA
    1085 50 mA started 19:52
    1086 40 mA
    1087 30 mA
    1088 20 mA 20:30
    1089 15 mA
    1090 10 mA at 21:05
    1091 5 mA
    1092 3 mA, end of current scan, 21:43
    prepare for elastic ed, d(e,e') in L arm, Kin. Q1_b40 in R arm
    First Physics run: 1093 Ela_D10, L arm at 14.50 deg., 2.7112 GeV/c
    1094 Ela_D11;
    note that neither runs shows much of an elastic peak, but an overwhelming quasi-elastic spectrum is present; we do 1 GeV^2 next to end this campaign
    1095 Ela_D14